Question 5 options:
The average monthly mortgage payment including principal and interest is $982 in the United States. If the standard deviation is approximately $180 and the mortgage payments are approximately normally distributed, find the probabilities.
a) The selected monthly payment is more than $1175.
P(X > 1175) =
b) The selected monthly payment is between $700 and $1000.
P(700 < X < 1000) =

Answer by Boreal(15235)

z=(x-mean)/sd
so z>(193/180)=1.072
that probability is 0.1418
-
This is z=-282/180 and z=18/180 for probability 0.4816
check with calculator 2ndVARS(700,1000,982,180)ENTER for 0.4812. The difference is that I rounded -282,180 and the calculator doesn't round until the end. Ask how much they want the z-value rounded to.